# Note about non-compact embeddings between Besov spaces

### v1

@@ 3. Embeddings Between Sequence Spaces
ii: By Corollary 2.17 i the map is noncompact. The finitely strictly singularity follows from Proposition 4.2 i.
v: When $p_{0}>p_{1}$, i.e. the case $s_{0}-s_{1}=\frac{n}{p_{0}}-\frac{n}{p_{1}}>0$ and $q_{0}=q_{1}$, the result follows from Proposition 4.2 ii. The only remaining case left from the previous discussion is the case $p_{0}=p_{1}$. Recall that either $s_{0}-s_{1}>0$, or $s_{0}-s_{1}=0$ but $q_{0}\leq q_{1}$.
@@ 5. On Homogeneous Besov Embeddings
Recall that the homogeneous Besov space ˙$\dot{\mathscr{B}}_{pq}^{s}(\mathbb{R}^{n})$ is defined to be the element in the quotient space $\dot{\mathscr{S}}^{\prime}(\mathbb{R}^{n})=\mathscr{S}^{\prime}(\mathbb{R}^{n})/\{\text{polynomials}\}$ for which
@@
Here $(\varphi_{j})_{j\in\mathbb{Z}}\subset\dot{\mathscr{S}}(\mathbb{R}^{n})$ is a fixed family of Schwartz functions with all moment vanishings, such that $\operatorname{supp}\hat{\varphi}_{0}\subset\{\frac{1}{2}<|\xi|<2\}$, $\hat{\varphi}_{j}(\xi)=\hat{\varphi}_{0}(2^{-j}\xi)$ for all $j\in\mathbb{Z}$, and $\sum_{j\in\mathbb{Z}}\hat{\varphi}_{j}(\xi)=0$ for all $\xi\neq 0$. See e.g. [Tri10, Chapter 5] for details.
@@
Let $p_{0},p_{1},q_{0},q_{1}\in(0,\infty]$, $s_{0},s_{1}\in\mathbb{R}$. Consider the embedding ˙$\dot{\mathscr{B}}_{p_{0}q_{0}}^{s_{0}}(\mathbb{R}^{n})\hookrightarrow\dot{\mathscr{B}}_{p_{1}q_{1}}^{s_{1}}(\mathbb{R}^{n})$.
@@
    ˙$\dot{\mathscr{B}}_{p_{0}q_{0}}^{s_{0}}(\mathbb{R}^{n})\not\subset\dot{\mathscr{B}}_{p_{1}q_{1}}^{s_{1}}(\mathbb{R}^{n})$, i.e. there is no embedding, if and only if one of the following occurs:
@@
If we adapt the notation of index sets in Proposition 2.12, the orthonormal basis in the proof below can be written as $\{\dot{\psi}_{jm}^{g}:j\in\mathbb{Z},m\in 2^{-j}\mathbb{Z}^{n},g\in\{0,1\}^{n}\backslash\{0\}^{n}\}\subset\dot{\mathscr{S}}(\mathbb{R}^{n})$. It is more convenient to use identifications $(2^{-j}\mathbb{Z}^{n})\times(\{0,1\}^{n}\backslash\{0\}^{n})\simeq\mathbb{N}_{0}$ on the index $(m,g)$, because we do not need the analogy of Proposition 2.12 ii in the proof.
@@
i: In this assumption $\ell^{q_{0}}(2^{j(s_{0}-\frac{n}{p_{0}})}\cdot\ell^{p_{0}})_{j=-\infty}^{\infty}\not\subset\ell^{q_{1}}(2^{j(s_{1}-\frac{n}{p_{1}})}\cdot\ell^{p_{1}})_{j=-\infty}^{\infty}$, thus ˙$\dot{\mathscr{B}}_{p_{0}q_{0}}^{s_{0}}(\mathbb{R}^{n})\not\subset\dot{\mathscr{B}}_{p_{1}q_{1}}^{s_{1}}(\mathbb{R}^{n})$.
@@
When $s_{0}-s_{1}=\frac{n}{p_{0}}-\frac{n}{p_{1}}=:s$, the map $J_{s}(x_{j})_{j\in\mathbb{Z}}:=(2^{js}x_{j})_{j\in\mathbb{Z}}$ defines isomorphisms that for $i=0,1$, $J_{s}:\ell^{q_{i}}(2^{js}\cdot\ell^{p_{i}})_{j\in\mathbb{Z}}\to\ell^{q_{i}}(\mathbb{Z};\ell^{p_{i}})$ Since $\ell^{q_{0}}(\ell^{p_{0}})\not\subset\ell^{q_{1}}(\ell^{p_{1}})$ holds when $q_{0}>q_{1}$ or $p_{0}>p_{1}$, pulling back from $J_{s}$ we get the result.
ii: In this assumption, applying Proposition 3.5 the map $\ell^{q_{0}}(\mathbb{Z};\ell^{p_{0}})\hookrightarrow\ell^{q_{1}}(\mathbb{Z};\ell^{p_{1}})$ is finitely strictly singular. Clearly it is non-compact. Pulling back from $J_{s}$, we get the non-compactness and finitely strictly singularity of $\ell^{q_{0}}(2^{j(s_{0}-\frac{n}{p_{0}})}\cdot\ell^{p_{0}})_{j=-\infty}^{\infty}\hookrightarrow\ell^{q_{1}}(2^{j(s_{1}-\frac{n}{p_{1}})}\cdot\ell^{p_{1}})_{j=-\infty}^{\infty}$, so is ˙$\dot{\mathscr{B}}_{p_{0}q_{0}}^{s_{0}}(\mathbb{R}^{n})\hookrightarrow\dot{\mathscr{B}}_{p_{1}q_{1}}^{s_{1}}(\mathbb{R}^{n})$.
iii: In this assumption, when $q_{0}=q_{1}$, the result follows from Lemma 3.1; when $p_{0}=p_{1}$ the result from the fact that for every $j\in\mathbb{Z}$, the spaces $2^{j(s_{0}-\frac{n}{p_{0}})}\cdot\ell^{p_{0}}=2^{j(s_{1}-\frac{n}{p_{1}})}\cdot\ell^{p_{1}}$ and have the same norms. ∎
@@ 6. Further Remarks
In Theorems 1.1 and 1.2, we observe the impact of differences between function spaces on bounded Lipschitz domains $\Omega$ and on $\mathbb{R}^{n}$. A natural question that arises is what results can be obtained for embeddings on unbounded open sets.
We believe that the upper bounds in Proposition 3.5 could be sharpened, and that it should be possible to derive corresponding lower bounds. This could lead to improved estimates on the rate of decay for Bernstein numbers between Besov spaces when the embedding is non-compact but finitely strictly singular. The case where $\Omega=\mathbb{R}^{n}$ is particularly intriguing.
Given the obtained results for Besov spaces, it seems natural to ask about the inner structure for non-compact embeddings between Triebel-Lizorkin spaces. This could be quite interesting, given their different behaviors of the corresponding sequence spaces from the wavelet decomposition, which makes the study of Triebel-Lizorkin spaces more challenging.

### v2

@@ 3. Embeddings Between Sequence Spaces
$i$: By Corollary 2.18 i the map is noncompact. The finitely strictly singularity follows from Proposition 4.2 i.
iii: When $p_{0}>p_{1}$, i.e. the case $s_{0}-s_{1}=\frac{n}{p_{0}}-\frac{n}{p_{1}}>0$ and $q_{0}=q_{1}$, the result follows from Proposition 4.2 ii. The only remaining case left from the previous discussion is the case $p_{0}=p_{1}$. Recall that either $s_{0}-s_{1}>0$, or $s_{0}-s_{1}=0$ but $q_{0}\leq q_{1}$.
@@ 5. On Homogeneous Besov Embeddings
Recall that the homogeneous Besov space $\dot{\mathscr{B}}_{pq}^{s}(\mathbb{R}^{n})$ is defined to be the element in the quotient space $\dot{\mathscr{S}}^{\prime}(\mathbb{R}^{n})=\mathscr{S}^{\prime}(\mathbb{R}^{n})/\{\text{polynomials}\}$ for which
@@
Here $(\varphi_{j})_{j\in\mathbb{Z}}\subset\dot{\mathscr{S}}(\mathbb{R}^{n})$ is a fixed family of Schwartz functions with all moment vanishings, such that $\operatorname{supp}\hat{\varphi}_{0}\subset\{\frac{1}{2}<|\xi|<2\}$, $\hat{\varphi}_{j}(\xi)=\hat{\varphi}_{0}(2^{-j}\xi)$ for all $j\in\mathbb{Z}$, and $\sum_{j\in\mathbb{Z}}\hat{\varphi}_{j}(\xi)=1$ for all $\xi\neq 0$. See e.g. [Tri10, Chapter 5] for details.
@@
Let $p_{0},p_{1},q_{0},q_{1}\in(0,\infty]$, $s_{0},s_{1}\in\mathbb{R}$. Consider the embedding $\dot{\mathscr{B}}_{p_{0}q_{0}}^{s_{0}}(\mathbb{R}^{n})\hookrightarrow\dot{\mathscr{B}}_{p_{1}q_{1}}^{s_{1}}(\mathbb{R}^{n})$.
@@
    $\dot{\mathscr{B}}_{p_{0}q_{0}}^{s_{0}}(\mathbb{R}^{n})\not\subset\dot{\mathscr{B}}_{p_{1}q_{1}}^{s_{1}}(\mathbb{R}^{n})$, i.e. there is no embedding, if and only if one of the following occurs:
@@
[…]
@@
i: In this assumption $\ell^{q_{0}}(2^{j(s_{0}-\frac{n}{p_{0}})}\cdot\ell^{p_{0}})_{j=-\infty}^{\infty}\not\subset\ell^{q_{1}}(2^{j(s_{1}-\frac{n}{p_{1}})}\cdot\ell^{p_{1}})_{j=-\infty}^{\infty}$, thus $\dot{\mathscr{B}}_{p_{0}q_{0}}^{s_{0}}(\mathbb{R}^{n})\not\subset\dot{\mathscr{B}}_{p_{1}q_{1}}^{s_{1}}(\mathbb{R}^{n})$.
@@
When $s_{0}-s_{1}=\frac{n}{p_{0}}-\frac{n}{p_{1}}=:s$, the map $J_{s}(x_{j})_{j\in\mathbb{Z}}:=(2^{js}x_{j})_{j\in\mathbb{Z}}$ defines isomorphisms that for $i=0,1$, $J_{s}:\ell^{q_{i}}(2^{js}\cdot\ell^{p_{i}})_{j\in\mathbb{Z}}\to\ell^{q_{i}}(\mathbb{Z};\ell^{p_{i}})$ . Since $\ell^{q_{0}}(\ell^{p_{0}})\not\subset\ell^{q_{1}}(\ell^{p_{1}})$ holds when $q_{0}>q_{1}$ or $p_{0}>p_{1}$, pulling back from $J_{s}$ we get the result.
ii: In this assumption, applying Proposition 3.5 the map $\ell^{q_{0}}(\mathbb{Z};\ell^{p_{0}})\hookrightarrow\ell^{q_{1}}(\mathbb{Z};\ell^{p_{1}})$ is finitely strictly singular. Clearly it is non-compact. Pulling back from $J_{s}$, we get the non-compactness and finitely strictly singularity of $\ell^{q_{0}}(2^{j(s_{0}-\frac{n}{p_{0}})}\cdot\ell^{p_{0}})_{j=-\infty}^{\infty}\hookrightarrow\ell^{q_{1}}(2^{j(s_{1}-\frac{n}{p_{1}})}\cdot\ell^{p_{1}})_{j=-\infty}^{\infty}$, so is $\dot{\mathscr{B}}_{p_{0}q_{0}}^{s_{0}}(\mathbb{R}^{n})\hookrightarrow\dot{\mathscr{B}}_{p_{1}q_{1}}^{s_{1}}(\mathbb{R}^{n})$.
iii: When $q_{0}=q_{1}$, the result follows from Lemma 3.1; when $p_{0}=p_{1}$ the result from the fact that for every $j\in\mathbb{Z}$, the spaces $2^{j(s_{0}-\frac{n}{p_{0}})}\cdot\ell^{p_{0}}=2^{j(s_{1}-\frac{n}{p_{1}})}\cdot\ell^{p_{1}}$ and have the same norms. ∎
@@ 6. Further Remarks
In Theorems 1.1 and 1.2, we observe the impact of differences between function spaces on bounded Lipschitz domains $\Omega$ and on $\mathbb{R}^{n}$. A natural question that arises is what results can be obtained for embeddings on unbounded open sets (say unbounded uniform domains).
We believe that the upper bounds in Proposition 3.5 could be sharpened, and that it should be possible to derive corresponding lower bounds. This could lead to improved estimates on the rate of decay for Bernstein numbers between Besov spaces when the embedding is non-compact but finitely strictly singular. Note that although we do not have a canonical definition of Besov norms, the decay rates of their Bernstein numbers are determined up to a constant multiple which still make sense. The case where $\Omega=\mathbb{R}^{n}$ is particularly intriguing, which is equivalent to computing the decay rate of $b_{n}\big(\ell^{q_{0}}(\ell^{p_{0}})\to\ell^{q_{1}}(\ell^{p_{1}})\big)$.
Given the obtained results for Besov spaces, it seems natural to ask about the inner structure for non-compact embeddings between Triebel-Lizorkin spaces. This could be quite interesting, given their different behaviors of the corresponding sequence spaces from the wavelet decomposition, which makes the study of Triebel-Lizorkin spaces more challenging. Also, it is worth noting that the obtained results for Besov spaces could be used for non-compact Sobolev embeddings which will generalize and extend results from [LM23] and [BG89].